OpenStudy (anonymous):

What is the range of f(x)=√(4-3x^2) and how do you find it?

OpenStudy (anonymous):

Can a square root be a negative number?

OpenStudy (anonymous):

yeah because i mean the square root of 4 is 2 and - 2 right?

OpenStudy (anonymous):

hm. Typically you say that the square root of x is positive... the range is zero and positive real numbers. But I have to admit I am temporarily confused, since you are right, (-2)^2 = 4

OpenStudy (anonymous):

well i feel like that can not have what is being squared be a negative because you can not take the square of a negative because that doesn't exist.

OpenStudy (anonymous):

The deal is that the domain for y = sqrt(x) is real numbers greater than or equal to zero. You can't put in a negative number as x and keep y as real numbers... So given that x must be real and >= 0, the range for y is also real numbers >= 0

OpenStudy (anonymous):

The range is the allowable results generated from the allowable inputs... the domain. So since x = -2 is not in the domain, you don't have to consider it... the x's must be >= 0, so the only resulting y's are also real and >=0

OpenStudy (anonymous):

back to your real problem... you should first find the domain by finding the x values that make the expression under the square root "legal"... that expression must be >=0.

OpenStudy (anonymous):

its not a problem of x being negative or not because a negative number squared is always positive. I think x has to be small enough so that you dont take the square of a negative number.

OpenStudy (anonymous):

yes :) in your real problem, that is the situation... I was sidetracked on the sqrt(4) issue. x must be small enough to avoid the negative under the root... and notice, x itself can be negative, as long as x^2 is still small enough.

OpenStudy (anonymous):

the domain (-∞, 4/3)

ganeshie8 (ganeshie8):

there is a nice alternative for finding range, here

OpenStudy (anonymous):

I don't think you can choose really large negative x's... makes the square root argument negative. maybe it's just between (-4/3, 4/3) ?

OpenStudy (anonymous):

@ganeshie8 was there a link? I'm totally interested... there has to be a better way than what I am doing!!

OpenStudy (anonymous):

@ganeshie8 i second that

ganeshie8 (ganeshie8):

im not sure of links.. mukushla taught me this few days back : lets say, \(\sqrt{4-3x^2} = k\) square both sides, \(4-3x^2 = k^2\) \(-3x^2+(4-k^2) = 0 \)

ganeshie8 (ganeshie8):

since we are looking for real x, \(b^2-4ac >= 0 \) => \(0-4(-3)(4-k^2) >= 0 \) \(-k^2 >= 4 \) \(k^2 <= 4\) \(k <= 2\)

ganeshie8 (ganeshie8):

since \(k\) cannot be negative, the range is [0, 2]

OpenStudy (anonymous):

interesting, it is...

OpenStudy (anonymous):

but how can it be 2 if that would be taking the swuare root of negative 12?

ganeshie8 (ganeshie8):

yea it helped me on many occasions with my problems of finding range

OpenStudy (anonymous):

[0,2] is the range... it's the range of y values, not the part under the sqrt

OpenStudy (anonymous):



OpenStudy (anonymous):

oh my gosh.... yeah you are so right i was totally comfised for a second i was thinking domain!!!